OpenStudy (anonymous):

(6+2)^2/16 + 3 x 9, is it 6.75, I'm just checking

OpenStudy (lalaly):

6+2=8 8^2=64 64/16=4 4+(3x9) = 4+27 = 31

OpenStudy (anonymous):

Ohhhh, that makes more sence

OpenStudy (lalaly):

=)